Find the general solution to the non-homogeneous system of DE: -4 X+
Answer #1

Ang -4-2 5 -4 4-2) giem w]+[ > salotion to homogeneous peut q- []a eigenvalues [A-91) =0 > | 64-2)(4->) +20=0 -16-42+42 +22+t + (6) 50 din 377 (-40, now particular solution: 2pct) = 0++ € de nt -4d, t-46, +5dat + 562 da. + -4dit 46, +4 det +462, -4d